OpenStudy (anonymous):

Write the standard equation of the circle with center (2, 3) and a diameter of 12.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Hi give me a min

OpenStudy (anonymous):

(x-2)^2+(y-3)^2=6 this is because the equation is (x-h)^2+(y-k)^2=r

OpenStudy (anonymous):

h is the x point

OpenStudy (anonymous):

k is the y point and the r is the raidus

OpenStudy (anonymous):

go it?

OpenStudy (anonymous):

yes thank you! can I ask you another question?

OpenStudy (anonymous):

sure make another post

OpenStudy (anonymous):

1. The equation of a circle is (x + 7)2 + (y + 2)2 = 49. Determine the length of the diameter. 2. The equation of a circle is (x + 3)2 + (y + 7)2 = 25. Where is (3, 4) located in relation to the circle? Answer A. On the circle B. In the interior of the circle C. In the exterior of the circle D. at the center of the circle I had 2 questions lol

OpenStudy (anonymous):

1. Since 49 is the radius to find the diameter it is just 49*2

OpenStudy (anonymous):

Number 1 only gives me the anwsers A. 7 B. 49 C. 9 D. 14

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

silly me the last part is r^2

OpenStudy (anonymous):

so it is √(49)*2

OpenStudy (anonymous):

so 7?

OpenStudy (anonymous):

also the first question is 36 not 6

OpenStudy (anonymous):

no since 7 is the radius

OpenStudy (anonymous):

the diameter is r*2 so it is 14

OpenStudy (anonymous):

also remember the first question that i did with you it is not 6 it is 36 for r

OpenStudy (anonymous):

okay thank you! what about the other question I asked?

OpenStudy (anonymous):

ok give me a second

OpenStudy (anonymous):

i was going to draw it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the a is the (3,4)

OpenStudy (anonymous):

so that would make it interior b/c its smaller then (3,7)?

OpenStudy (anonymous):

the circle is negative because the circle equation is (x-h)^2+(y-h)^2=r^2 not (x+h)^2+(y+h)^2=r^2

OpenStudy (anonymous):

then exterior?

OpenStudy (anonymous):

no it would be exterior since it is outside of the circle

OpenStudy (anonymous):

ohh i get

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ok gtg have a good day

OpenStudy (anonymous):

Np